OpenStudy (anonymous):

Find each function value, it if exists. f(t)=sqrt t^2+1 f(0) = ?

OpenStudy (anonymous):

hello bunny

OpenStudy (anonymous):

Hi satellite. Doing square roots and I am ill...figuratively LOL

OpenStudy (anonymous):

replace t by 0 to get \[f(0)=\sqrt{0^2+1}=\sqrt{0+1}=\sqrt{1}=1\]

OpenStudy (anonymous):

i assume you meant \[f(t)=\sqrt{t^2+1}\] and not \[f(t)=\sqrt{t^2}+1\]

OpenStudy (anonymous):

yes...now what if it is f(-1)